### v2

@@ Abstract
Weakly-supervised segmentation (WSS) has emerged as a solution to mitigate the conflict between annotation cost and model performance by adopting sparse annotation formats (e.g., point, scribble, block, etc.). Typical approaches attempt to exploit anatomy and topology priors to directly expand sparse annotations into pseudo-labels. However, due to lack of attention to the ambiguous edges in medical images and insufficient exploration of sparse supervision, existing approaches tend to generate erroneous and overconfident pseudo proposals in noisy regions, leading to cumulative model error and performance degradation. In this work, we propose a novel WSS approach, named ProCNS, encompassing two synergistic modules devised with the principles of progressive prototype calibration and noise suppression. Specifically, we design a Prototype-based Regional Spatial Affinity (PRSA) loss to maximize the pair-wise affinities between spatial and semantic elements, providing our model of interest with more reliable guidance. The affinities are derived from the input images and the prototype-refined predictions. Meanwhile, we propose an Adaptive Noise Perception and Masking (ANPM) module to obtain more enriched and representative prototype representations, which adaptively identifies and masks noisy regions within the pseudo proposals, reducing potential erroneous interference during prototype computation. Furthermore, we generate specialized soft pseudo-labels for the noisy regions identified by ANPM, providing supplementary supervision. Extensive experiments on three medical image segmentation tasks involving different modalities demonstrate that the proposed framework significantly outperforms representative state-of-the-art methods. Code and data are available at https://github.com/LyxDLiI/ProCNS.git.
@@ 1 Introduction
Weakly-supervised segmentation (WSS) has emerged as a promising solution by employing sparse annotations, such as points, scribbles, blocks, and others (as illustrated in the top panel of Fig. 1), to train segmentation models, effectively alleviating the inherent conflict between annotation cost and model performance. Existing methods can be mainly categorized into pseudo-proposal, consistency learning, and auxiliary task methods. Pseudo-proposal methods [3, 4, 5, 6] employ prior knowledge, semantic affinity or model prediction to expand and generate pseudo-labels from the original sparse annotations. They typically involve multi-stage training and are susceptible to noise accumulation. Consistency learning methods [7, 8, 9] penalize inconsistent predictions on different views of the same image to regularize the training process, yet they fail to exploit the semantic correlation between annotated and unannotated regions. Auxiliary task methods [10, 11] impose comprehensive constraints by incorporating additional tasks such as boundary prediction, which may nevertheless impair the performance of the main segmentation task.
@@
Figure 1: Top: Examples of an optical coherence tomography angiography (OCTA) image, a fundus image, and an endoscope image, coupled with their respective sparse annotations of diverse types. UA, BG, FAZ, OD, OC and PL respectively represent unlabeled region, background, foveal avascular zone, optic disc, optic cup and polyp. Bottom: Visualization of pseudo-label error maps generated by TreeEnergy [4], DMPLS [12] and our ProCNS.
@@
Among those aforementioned methods, pseudo-proposal approaches are most prevalent. Yet, a potentially overlooked pivotal detail is that when sparse labels are generated through either manual annotation or automated algorithms, e.g., Random Walks [13], object regions for annotation tend to be preferentially positioned within readily distinguishable regions (for example, the central regions of the foveal avascular zone and polyp), rather than nebulous and uncertain regions (for example, the boundary intersection regions between the optic disc and the optic cup). Intuitively, those pseudo-labels predominantly inhabit less-informative regions rather than hard-yet-informative ones. The former, easily classified by a model even under the supervision of sparse labels, sharply contrasts the latter; it often exhibits significant prediction fluctuation and unreliability throughout the training process. The skewed annotation proportion, favoring less-informative regions (often the majority) over their hard-yet-informative counterparts (often the minority), may be detrimental to model training. Specifically, under the supervision of such sparse labels, the trained models exhibit a tendency to allocate predictions more extensively to less-informative regions. The diminutive and steady loss values observed in less-informative regions and the pixel-wise averaging characteristic of segmentation losses, e.g., the partial Cross-Entropy (pCE) loss [14], diminish the efficacy of hard-yet-informative regions, subsequently leading to erroneous predictions at the boundary regions (as illustrated in the bottom panel of Fig. 1). In medical images, the structures and lesions tend to be inherently more ambiguous than those in natural images, exacerbating the aforementioned issue. Direct or indirect utilizations of those erroneous predictions as pseudo-labels may induce further error accumulation, leading to performance degradation.
@@
The most direct solution to the above issue is to increase the coverage and proportion of annotations for hard-yet-informative regions, yet it conflicts with the objective of reducing manual annotation costs in WSS. Consequently, it is natural to propose prototype representation learning to address the issue. Prototype representation learning has been explored and validated in few-shot and semi-supervised learning tasks [15, 16, 17], which can effectively summarize class representations and generate reliable pseudo-labels. However, in the context of WSS, prototypes extracted from sparse annotations or noisy pseudo-labels lack semantic richness and sufficient accuracy. Inaccurate prototypes may result in misclassifications of unannotated regions. Adaptive approaches that perceive and mask noisy regions while utilizing as many unannotated regions from diverse target classes as possible to generate prototypes could potentially alleviate this issue. However, such explorations are relatively rare.
@@
In such context, we propose a novel weakly-supervised medical image segmentation algorithm, named ProCNS, encompassing two complementary modules conforming to the principles of progressive prototype representation refinement and noise suppression. Firstly, we formulate a Prototype-based Regional Spatial Affinity (PRSA) loss to maximize both spatial and semantic pair-wise affinities, thereby providing our model of interest with more robust guidance. The affinities are extracted from the input images and the prototype-refined predictions. Simultaneously, an Adaptive Noise Perception and Masking (ANPM) module is designed to progressively identify and mask noisy regions within the pseudo proposals, mitigating the risk of erroneous interference during prototype computation. In addition, the prototype-refined predictions are harnessed to generate soft pseudo-labels specialized for noisy regions identified by ANPM, providing additional supervision. The main contributions of this paper are summarized as follows:
@@
  • To the best of our knowledge, this is the first attempt to employ progressive prototype calibration and noise suppression to address the insufficiency of prototype semantic representativeness and richness in WSS. Moreover, the proposed ProCNS can be flexibly utilized as a seamless integration plugin for existing WSS methods.
@@
  • We evaluate ProCNS on three different medical image segmentation tasks involving various forms of sparse annotations. Experimental results showcase the superiority of ProCNS over existing comparative methods.
@@ 2.1 Weakly-supervised Segmentation
Weakly-supervised segmentation aims to reduce the annotation cost by training segmentation models on data with inexact annotations. Existing methods fall into three main categories: pseudo-proposal, consistency learning and auxiliary task methods. Pseudo-proposal methods [4, 18, 19, 20] employ prior knowledge, semantic similarity or model predictions to propagate sparse labels to unlabeled regions, thereby extending and generating pseudo-labels. For instance, [4] employ the minimum spanning tree property to design a tree filter for effectively mitigating pseudo-labels’ noise. Compete-to-win [3] compares multiple confidence maps produced by auxiliary branches to vote for the best one to serve as the pseudo-label. Consistency learning methods [7, 8, 9] generally utilize cross-view consistency to penalize inconsistent segmentation. For example, [7] adopt a mixup strategy to obtain images with diverse views, followed by a consistency loss to regularize the model training process. Auxiliary task methods [21, 10] enhance comprehensive constraints by integrating other tasks, such as sub-category exploration [21] and multi-label image classification [10].
@@ 2.2 Prototype Representation Learning
Given its capacity of clustering similar units into a unified embedding space, prototype representation learning can effectively capture the structures and features of data. It has been comprehensively studied and validated in few-shot, semi-supervised and unsupervised tasks [22, 23, 24, 25]. The effectiveness of utilizing prototypes to refine predictions has already been preliminarily explored. For instance, [26] introduce a multi-prototype classifier to replace the traditional parameterized classifier for generating predictions, while [27] utilize sample-wise prototypes to generate cross-sample probability predictions. However, in the WSS setting, the target prototypes generated from sparse labels may lack semantic representativeness. Directly or indirectly utilizing them for prediction refinement could potentially result in overconfident errors. The idea of adaptively selecting regions with less noise to generate prototypes holds the promise of mitigating this issue. Currently, such explorations are relatively rare and our work aims to fill this research gap.
@@ 3.3.3 Affinity Loss
Our affinity loss is designed to facilitate the propagation of region-level semantic information from reliable regions to noisy ones by maximizing the affinity between pixels. Feeding an image to the model, the model outputs the logits $p$.
Following [29], we utilize a Gaussian kernel function to design the low-level weight function $\omega^{\rm low}$, which is defined by the distinction between two pixels in terms of image intensity value $v$ and spatial location $l$
@@ 3.6 Total Objective Formulation
Following [31] and [7], we employ the pCE loss [14] to provide direct supervision by matching predictions with sparse labels, which is expressed as
@@
Significantly, the final model in the Main stage is derived by extending the training of the model from the Initialization stage, where the training epochs for the former approximately account for one-tenth of the total training epochs. Furthermore, ANPM achieves prototype calibration by progressively replacing the labels $y$ used for prototype computation in Eq. 3 (i.e., replacing $\hat{y}^{\rm d}_{t-1}$ with $\hat{y}^{\rm d}_{t}$). This ensures that network architectures and the quantity of optimized parameters remain constant throughout the entire training process. Thus, ProCNS can be regarded as an approximate end-to-end framework that joints two internal training paradigms.
Table 1: Summary of the three sparsely-annotated datasets. The proportion denotes the average percentage of the sparsely-annotated regions over the fully-annotated ones. The relative time cost (RTC) denotes the approximate percentage of the average time consumed in manually annotating images with sparse and full annotations.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate our approach on the sOCTA [32], RIM-ONE [33] and Kvarsir-SEG [34] datasets, respectively utilizing point, scribble and block annotations for the FAZ, ODOC and polyp segmentation tasks (as delineated in Table 1). The sparse annotations are generated using different automated algorithms based on the original full annotations. The sOCTA, RIM-ONE and Kvarsir-SEG datasets consist of 708, 99, and 900 training samples, as well as 304, 60, and 100 testing samples. For all three datasets, we randomly split 20% of the training samples for validation. For data pre-processing, the fundus images, OCTA images and endoscope images are respectively resized to $384\times 384$, $256\times 256$ and $384\times 384$ as network inputs. Then, we normalize the intensity values of each image to have zero mean and unit variance. Additionally, random rotation and flipping are applied for data augmentation.
@@ 4.2 Implementation Details
All compared WSS methods and ProCNS are implemented with PyTorch on Nvidia RTX 3090 GPUs. We employ the vanilla UNet [35] as the network architecture, with its encoder and decoder respectively comprising four downsampling blocks and four upsampling blocks. The highest-dimensional embedding is 256. The SGD optimizer with a momentum of $0.9$ and a weight decay of $1e^{-4}$ is adopted. The polynomial decay schedule is employed with the initial learning rate for each task being $1\times 10^{-2}$. The EMA decay rate is set to $0.8$. The trade-off coefficients $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are respectively set to $0.1,0.5,0.1$ and $0.01$. The $\sigma_{l}$, $\sigma_{v}$ and radius $r$ of the neighborhood window are respectively set to 6, 0.1 and 5. The number of training epochs during the Initialization stage is respectively set to $100,400$ and $100$ for each of the three segmentation tasks.
We now provide detailed discourse regarding the acquisition of the three forms of sparse annotations in the following text. The generation of sparse annotations is facilitated by three automated algorithms, which are implemented utilizing OpenCV2, SimpleITK and Scikit-image. For points, inspired by [36] and [37], we automatically generate points from full annotations. Specifically, we extract the maximum inscribed rectangles from the objects. The four sides of each rectangle are contracted inward by a predetermined value. We then obtain the midpoints of the diminished rectangle’s sides. These midpoints, encompassing multiple pixels, are then convolved with a discrete 2D Gaussian kernel to simulate a manual brushstroke. For scribbles, following [12], a cross-shaped kernel is utilized to perform morphological erosion from the center pixel of each full annotation (through the erode function in OpenCV2). Subsequently, the residual regions of disparate classes are compressed into thin lines (through the skeletonize function in Scikit-image), thus generating centerlines or skeletons as scribble annotations. Following [4], block annotations are derived via morphological erosion transformations starting from the edges and progressing towards the center.

### v3

@@ Abstract
Weakly-supervised segmentation (WSS) has emerged as a solution to mitigate the conflict between annotation cost and model performance by adopting sparse annotation formats (e.g., point, scribble, block, etc.). Typical approaches attempt to exploit anatomy and topology priors to directly expand sparse annotations into pseudo-labels. However, due to lack of attention to the ambiguous boundaries in medical images and insufficient exploration of sparse supervision, existing approaches tend to generate erroneous and overconfident pseudo proposals in noisy regions, leading to cumulative model error and performance degradation. In this work, we propose a novel WSS approach, named ProCNS, encompassing two synergistic modules devised with the principles of progressive prototype calibration and noise suppression. Specifically, we design a Prototype-based Regional Spatial Affinity (PRSA) loss to maximize the pair-wise affinities between spatial and semantic elements, providing our model of interest with more reliable guidance. The affinities are derived from the input images and the prototype-refined predictions. Meanwhile, we propose an Adaptive Noise Perception and Masking (ANPM) module to obtain more enriched and representative prototype representations, which adaptively identifies and masks noisy regions within the pseudo proposals, reducing potential erroneous interference during prototype computation. Furthermore, we generate specialized soft pseudo-labels for the noisy regions identified by ANPM, providing supplementary supervision. Extensive experiments on six medical image segmentation tasks involving different modalities demonstrate that the proposed framework significantly outperforms representative state-of-the-art methods. Code and data are available at https://github.com/LyxDLiI/ProCNS.
@@ 1 Introduction
Weakly-supervised segmentation (WSS) has emerged as a promising solution by employing sparse annotations, such as points, scribbles, blocks and others (as illustrated in the top panel of Fig. 1), to train segmentation models, effectively alleviating the inherent conflict between annotation cost and model performance. Existing methods can be mainly categorized into pseudo-proposal, consistency learning, auxiliary task and distillation-based methods. Pseudo-proposal methods [3, 4, 5, 6] employ prior knowledge, semantic affinity or model prediction to expand and generate pseudo-labels from the original sparse annotations. They typically involve multi-stage training and are susceptible to noise accumulation. Consistency learning methods [7, 8] penalize inconsistent predictions on different views of the same image to regularize the training process, yet they fail to exploit the semantic correlation between annotated and unannotated regions. Auxiliary task methods [9, 10] impose comprehensive constraints by incorporating additional tasks such as boundary prediction, which may nevertheless impair the performance of the main segmentation task. Distillation-based methods [11, 12] employ teacher models to distill richer knowledge from sparse annotations, transferring it to student models. They inevitably increase model complexity and computational burden.
@@
Figure 1: Top: Examples of an optical coherence tomography angiography (OCTA) image, a fundus image, an endoscope image, a hematoxylin and eosin (H&E)-stained tissue image, a cardiac magnetic resonance image (Cardiac MRI) and a brain tumor magnetic resonance image (Brain Tumor MRI), coupled with their respective sparse annotations of diverse types. UA, BG, FAZ, OD, OC, PL, NC, LV, MYO, RV and BT respectively represent unlabeled region, background, foveal avascular zone, optic disc, optic cup, polyp, nuclei, left ventricle, myocardium, right ventricle and brain tumor. Bottom: Visualization of pseudo-label error maps generated by TreeEnergy [4], DMPLS [13] and our ProCNS.
@@
Among the aforementioned methods, pseudo-proposal approaches are most prevalent. Yet a potentially overlooked pivotal detail is that when sparse labels are generated through either manual annotation or automated algorithms, e.g., Random Walks [14], regions selected for annotation tend to be preferentially positioned within readily distinguishable regions (for example, the central regions of the foveal avascular zone and polyp), rather than nebulous and uncertain regions (for example, the boundary intersection regions between the optic disc and the optic cup). Intuitively, those pseudo-labels predominantly inhabit less-informative regions rather than hard-yet-informative ones. The former, easily classified by a model even under the supervision of sparse labels, sharply contrasts the latter; it often exhibits significant prediction fluctuation and unreliability throughout the training process. The skewed annotation proportion, favoring less-informative regions (often the majority) over their hard-yet-informative counterparts (often the minority), may be detrimental to model training. Specifically, under the supervision of such sparse labels, the trained models exhibit a tendency to allocate predictions more extensively to less-informative regions. The diminutive and steady loss values observed in less-informative regions and the pixel-wise averaging characteristic of segmentation losses, e.g., the partial Cross-Entropy (pCE) loss [15], diminish the efficacy of hard-yet-informative regions, subsequently leading to erroneous predictions at the boundary regions (as illustrated in the bottom panel of Fig. 1). In medical images, the structures and lesions tend to be inherently more ambiguous than those in natural images, exacerbating the aforementioned issue. Direct or indirect utilization of those erroneous predictions as pseudo-labels may induce further error accumulation, leading to performance degradation.
@@
The most direct solution to the above issue is to increase the coverage and proportion of annotations for hard-yet-informative regions, which, however, conflicts with the objective of reducing manual annotation costs in WSS. Consequently, it is natural to propose prototype representation learning to address the issue. Prototype representation learning has been explored and validated in few-shot and semi-supervised learning tasks [16, 17, 18], which can effectively summarize class representations and generate reliable pseudo-labels. However, in the context of WSS, prototypes extracted from sparse annotations or noisy pseudo-labels lack semantic richness and sufficient accuracy. Inaccurate prototypes may result in misclassifications of unannotated regions. Adaptive approaches that perceive and mask noisy regions while utilizing as many unannotated regions from diverse target classes as possible to generate prototypes could alleviate this issue. However, such explorations are relatively rare.
@@
In such context, we propose a novel weakly-supervised medical image segmentation algorithm, named ProCNS, encompassing two complementary modules conforming to the principles of progressive prototype representation refinement and noise suppression. Firstly, we formulate a Prototype-based Regional Spatial Affinity (PRSA) loss to maximize spatial and semantic pair-wise affinities, thereby providing our model of interest with more robust guidance. The affinities are extracted from the input images and the prototype-refined predictions. Simultaneously, an Adaptive Noise Perception and Masking (ANPM) module is designed to progressively identify and mask noisy regions within the pseudo proposals, mitigating the risk of erroneous interference during prototype computation. In addition, the prototype-refined predictions are harnessed to generate soft pseudo-labels for the noisy regions identified by ANPM, providing additional supervision. The main contributions of this paper are summarized as follows:
@@
  • To the best of our knowledge, this work is the first attempt to employ progressive prototype calibration and noise suppression to address the insufficiency of prototype semantic representativeness and richness in WSS. Moreover, the proposed ProCNS can be flexibly utilized as a seamless integration plugin for existing WSS methods.
@@
  • We evaluate ProCNS on six different medical image segmentation tasks involving various forms of sparse annotations. Experimental results showcase the superiority of ProCNS over existing comparative methods.
@@ 2.1 Weakly-supervised Segmentation
Weakly-supervised segmentation aims to reduce the annotation cost by training segmentation models on data with inexact annotations. Existing methods fall into four main categories: pseudo-proposal, consistency learning, auxiliary task and distillation-based methods. Pseudo-proposal methods [4, 19, 20, 21] employ prior knowledge, semantic similarity or model predictions to propagate sparse labels to unlabeled regions, thereby generating extended pseudo-labels. For instance, Liang et al. [4] employ the minimum spanning tree property to design a tree filter for effectively mitigating pseudo-labels’ noise. Compete-to-win [3] compares multiple confidence maps produced by auxiliary branches to vote for the best one to serve as the pseudo-label. Consistency learning methods [7, 8] generally utilize cross-view consistency to penalize inconsistent segmentation. For example, Zhang et al. [7] adopt a mixup strategy to obtain images with diverse views, followed by a consistency loss to regularize the model training process. Auxiliary task methods [22, 9, 23] enhance comprehensive constraints by integrating other tasks, such as sub-category exploration [22] and multi-label image classification [9]. Additionally, Yang et al. [23] propose an innovative self-supervised auxiliary task based on contrastive learning to facilitate the downstream segmentation task. Distillation-based methods [11, 24, 12, 25] employ teacher models to distill more richer knowledge from sparse annotations, subsequently transferring to student models. For instance, in the context of WSS, Zhang et al. [24] develop a self-dual teaching architecture that leverages two-fold information cues, namely the discriminative object region and the full object region, to generate high-quality pseudo-labels, thereby better guiding the training of the student model.
@@ 2.2 Prototype Representation Learning
Given its capacity of clustering similar units into a unified embedding space, prototype representation learning can effectively capture the structures and features of data. It has been comprehensively studied and validated in few-shot, semi-supervised and unsupervised tasks [26, 27, 28, 29, 30, 31, 32]. The effectiveness of utilizing prototypes to refine predictions has already been preliminarily explored. For instance, Xu et al. [33] introduce a multi-prototype classifier to replace the traditional parameterized classifier, while Zhang et al. [34] utilize sample-wise prototypes to generate cross-sample probability predictions. However, in the WSS setting, the target prototypes generated from sparse labels may lack semantic richness. Directly or indirectly utilizing them for prediction refinement could potentially result in overconfident errors. The idea of adaptively selecting regions with less noise to generate prototypes holds the promise of mitigating this issue. Currently, such explorations are relatively rare and our work aims to fill this research gap.
@@ 3.3.3 Affinity Loss
Our affinity loss is designed to facilitate the propagation of region-level semantic information from reliable regions to noisy ones by maximizing the affinity between pixels. Feeding the image $x$ to the model, the model outputs the logits $p$.
Following Zhang et al. [36], we utilize a Gaussian kernel function to design the low-level weight function $\omega^{\rm low}$, which is defined by the distinction between two pixels within the image $x$ in terms of image intensity value $v$ and spatial location $l$,
@@ 3.6 Total Objective Formulation
Following Lee et al. [38] and Zhang et al. [7], we employ the pCE loss [15] to provide direct supervision by matching predictions with sparse labels, which is expressed as
@@
Notably, the Main stage’s model is derived by further training the model from the Initialization stage, wherein the training epochs for the latter approximately account for one-tenth of the total training epochs. Furthermore, ANPM achieves prototype calibration by progressively replacing the labels $y$ used for prototype computation in Eq. 3 (i.e., replacing $\hat{y}^{\rm d}_{t-1}$ with $\hat{y}^{\rm d}_{t}$ at epoch $t$), rather than introducing extra training parameters. Collectively, these points indicate that the network architectures and the quantity of the to-be-optimized parameters remain constant throughout the entire training process. Thus, ProCNS can be treated as an end-to-end framework that inter-connects two training paradigms.
Table 1: Summary of the six sparsely-annotated datasets. The proportion denotes the average percentage of the sparsely-annotated regions over the fully-annotated ones. The relative time cost (RTC) denotes the approximate percentage of the average time consumed in manually annotating images with sparse and full annotations.
@@
H&E WO Point 0.15% $\sim$2%
Cardiac MRI ACDC Scribble 11.03% $\sim$19%
Brain Tumor MRI BraTS2019 Block 64.70% $\sim$24%
@@ 4.1 Datasets and Evaluation Metrics
We evaluate our approach on the sOCTA [39], RIM-ONE [40], Kvarsir-SEG [41], WO [42], ACDC [43] and BraTS2019 [44] datasets, utilizing point, scribble and block annotations for the FAZ, ODOC, polyp, nuclei, cardiac multi-structures (CM) and whole brain tumor (WT) segmentation tasks (as delineated in Table 1). Among these six tasks, the cardiac multi-structures segmentation task employs manual scribble annotations provided by Valvano et al. [45], while all others employ sparse annotations generated using different automated algorithms based on their original full annotations.
The sOCTA, RIM-ONE, Kvarsir-SEG and WO datasets consist of 708, 99, 900 and 16 2D training samples, as well as 304, 60, 100 and 8 2D testing samples. The ACDC dataset consists of 80 3D training samples and 20 3D testing samples. For these five datasets, we randomly split 20% of the training samples for validation. The BraTS2019 dataset consists of 335 3D samples, each containing four modalities: FLAIR, T1, T1ce and T2. Following Luo et al. [46], we perform weakly-supervised whole brain tumor segmentation using only FLAIR images. The BraTS2019 samples are randomly split into 250 for training, 25 for validation and 60 for testing.
For data pre-processing, the OCTA images, fundus images and endoscope images are respectively resized to $256\times 256$, $384\times 384$ and $384\times 384$ as inputs. Following Yao et al. [47] and Qu et al. [48], the H&E-stained tissue images are first segmented into $250\times 250$ patches from the original size of $1000\times 1000$. The patches are then resized to $1024\times 1024$ using bicubic interpolation to serve as inputs. Following previous works [13] [49] [50], we convert the 3D volumes of cardiac MRI and brain tumor MRI into 2D slices (slices not containing the corresponding target are excluded) and the 2D slices are respectively resized to $256\times 256$ and $192\times 192$ as inputs. Then, we normalize the intensity values of all images to have zero mean and unit variance. Additionally, random rotation and flipping are applied for data augmentation [50].
@@ 4.2 Implementation Details
All compared WSS methods and ProCNS are implemented with PyTorch on Nvidia RTX 3090 GPUs. We employ the vanilla UNet [51] as the network architecture, with its encoder and decoder respectively comprising four downsampling blocks and four upsampling blocks. The highest-dimensional embedding is 256. The SGD optimizer with a momentum of $0.9$ and a weight decay of $1e^{-4}$ is adopted. The polynomial decay schedule is employed. The initial learning rate for the FAZ, ODOC, polyp tasks is uniformly set to $1\times 10^{-2}$, while for the nuclei, CM and WT tasks, it is respectively set to $5\times 10^{-2}$, $5\times 10^{-2}$ and $1\times 10^{-1}$. The EMA decay rate is set to $0.8$. The trade-off coefficients $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are respectively set to $0.1,0.5,0.1$ and $0.01$. The $\sigma_{l}$, $\sigma_{v}$ and radius $r$ of the neighborhood window are respectively set to 6, 0.1 and 5. The number of the training epochs during the Initialization stage is respectively set to 100, 400, 100, 5, 40 and 25 for the FAZ, ODOC, polyp, nuclei, CM and WT segmentation tasks.
We now provide detailed discourse regarding the acquisition of the three forms of sparse annotations in the following text. The generation of sparse annotations is facilitated by three automated algorithms, which are implemented utilizing OpenCV2, SimpleITK and Scikit-image. For points, inspired by Kim et al. [52], Zhang et al. [53] and Yao et al. [47], we automatically generate points from full annotations. Specifically, we extract the maximum inscribed rectangles from the objects. The four sides of each rectangle are contracted inward by a predetermined value. For the FAZ task, we extract the midpoints of the four sides, whereas for the nuclei task, we extract the rectangle’s center. These points, encompassing multiple pixels, are then convolved with a discrete 2D Gaussian kernel to simulate a manual brushstroke. For scribbles, following Luo et al. [13], a cross-shaped kernel is utilized to perform morphological erosion from the center pixel of each full annotation (through the erode function in OpenCV2). Subsequently, the residual regions of disparate classes are compressed into thin lines (through the skeletonize function in Scikit-image), thus generating centerlines or skeletons as scribble annotations. Following Liang et al. [4], block annotations are derived via morphological erosion transformations starting from the edges and progressing towards the center.